# Balancing the Past and Present: A Coordinated Replay Framework for Federated Class-Incremental Learning

### v1

@@ Abstract.
Federated learning, Consistent representation learning, Imbalanced data learning
ccs: Computing methodologies Distributed algorithms
1. Introduction
@@
Federated learning (FL) is a distributed machine learning paradigm that enables collaborative training of a shared global model across multiple data sources (McMahan et al., 2017; Hu et al., 2024a; Wang et al., [n. d.]; Fan et al., 2025). It periodically performs parameter-level interaction between clients and the server instead of gathering clients’ data, which can enhance data privacy while leveraging the diversity of distributed data sources to build a more generalized global model (Qi et al., 2023a; Fu et al., 2025; Hu et al., 2024b; Kairouz et al., 2021). Building upon this foundation, Federated Class-Incremental Learning (FCIL) extends FL by introducing dynamic data streams where clients sequentially encounter different task classes under non-independent and identically distributed data (Dong et al., 2022; Wu et al., 2025; Tran et al., 2024; Yang et al., 2024). However, this amplifies the inherent complexities of FL, as the global model must integrate heterogeneous and evolving knowledge from clients while mitigating catastrophic forgetting, despite having no or only limited access to historical data (Lu et al., 2024; Chen et al., 2024; Gao et al., 2024).
@@
Figure 1. Motivation of the FedCBDR. Traditional data replay strategies typically focus on local information and, due to the lack of global awareness, often result in imbalanced class distributions during replay. FedCBDR aims to explore global information in a privacy-preserving manner and leverage it for sampling, which can alleviate the class imbalance problem.
@@
To address the challenge of catastrophic forgetting in FCIL, data replay has emerged as a promising strategy for retaining knowledge from previous tasks. Existing replay-based methods can be broadly categorized into two types: generative-based replay and exemplar-based replay. The former leverages generative models to synthesize representative samples from historical tasks (Chen et al., 2024; Babakniya et al., 2023; Wang et al., 2024). Its core idea is to learn the data distribution of previous tasks and internalize knowledge in the form of model parameters, enabling the indirect reconstruction of prior knowledge through sample generation when needed (Tran et al., 2024; Zhang et al., 2023). However, they often overlook the computational cost of training generative models and are inherently constrained by the quality and fidelity of the synthesized data (Chen et al., 2024; Liang et al., 2024; Yoo and Park, 2024). In contrast, exemplar-based replay methods directly store real samples from previous tasks, avoiding the complexity of generative processes while leveraging high-quality raw data to ensure robust retention of prior knowledge (Dong et al., 2022; Liang et al., 2024; Li et al., 2024, 2025). These methods rely on a limited set of historical samples to maintain the decision boundaries of previously learned task classes. However, due to the lack of a global perspective on data distribution across clients, these methods are prone to class-level imbalance in replayed samples, which undermines the model’s ability to retain prior knowledge (Li et al., 2024, 2025).
@@
2. Related Work
2.1. Exemplar-based Replay Methods
@@
In FCIL, exemplar-based replay methods aim to mitigate catastrophic forgetting by storing and replaying a subset of samples from previous tasks. They typically maintain a small exemplar buffer on each client, which is used during training alongside new task data to preserve knowledge of previously learned classes (Dong et al., 2022; Li et al., 2024, 2025; Rebuffi et al., 2017). For example, GLFC alleviates forgetting in FCIL by leveraging local exemplar buffers for rehearsal, while introducing class-aware gradient compensation and prototype-guided global coordination to jointly address local and global forgetting (Dong et al., 2022). Moreover, Re-Fed introduces a Personalized Informative Model to strategically identify and replay task-relevant local samples, enhancing the efficiency of buffer usage and further reducing forgetting in heterogeneous client environments (Li et al., 2024). However, the lack of global insight in local sample selection often results in class imbalance, while the long-tailed distribution between replayed and current data is frequently overlooked, ultimately degrading the effectiveness of data replay (Li et al., 2024, 2025).
2.2. Generative-based Replay Methods
@@
Generative replay methods aim to reconstruct the samples of past tasks through techniques such as generative modeling (Tran et al., 2024; Chen et al., 2024; Zhang et al., 2023; Nguyen et al., 2024; Qi et al., 2023b), which enables the model to revisit historical knowledge to mitigate catastrophic forgetting. Following this line of thought, TARGET generates pseudo features through a globally pre-trained encoder and performs knowledge distillation by aligning the current model’s predictions with those of a frozen global model (Zhang et al., 2023); LANDER utilizes pre-trained semantic text embeddings as anchors to synthesize meaningful pseudo samples, and distills knowledge by aligning the model’s predictions with class prototypes derived from textual descriptions (Tran et al., 2024). However, these methods are typically limited by the high computational cost of training generative models and the suboptimal performance caused by low-fidelity pseudo samples (Tran et al., 2024; Zhang et al., 2023).
2.3. Knowledge Distillation-based Methods
@@
Knowledge distillation-based methods generally follow two paradigms. The first focus om aligning the output predictions of the current model with those of previous models, which aims to preserve task-specific decision boundaries (Li and Hoiem, 2017; Tan et al., 2024; Psaltis et al., 2023; Chen et al., 2025; Zhong et al., 2025). The second estimates the importance of model parameters for previously learned tasks and performs regularization to prevent forgetting (Kirkpatrick et al., 2017; Yu et al., 2024). Both approaches avoid storing raw data but are prone to knowledge degradation over time, especially as the number of tasks increases (Li and Hoiem, 2017; Kirkpatrick et al., 2017).
3. Preliminaries
@@
4. Class-wise Balancing Data Replay for Federated Class-Incremental Learning
This section presents an effective active data selection method for FCIL, which aims to explore global data distribution to balance class-wise sampling. Moreover, it leverages temperature scaling to adjust the logits, which can alleviate the imbalance between samples from previously learned and newly introduced tasks. Figure 2 illustrates the framework of the proposed method FedCBDR.
4.1. Global-perspective Data Replay (GDR)
@@
Figure 2. Illustration of the FedCBDR framework. It first trains local models using samples from current and previous tasks. After a fixed number of communication rounds, each client extracts local sample features using the global model and applies inverse singular value decomposition (ISVD) to obtain pseudo features. The server then aggregates both local models and pseudo features, performs SVD on the features, and selects representative samples based on leverage scores. The corresponding sample indices are sent back to the clients for balanced replay.
Inspired by Singular Value Decomposition (SVD) (Chai et al., 2024, 2022), we first generate a set of random orthogonal matrices: a client-specific matrix $P_{k}^{(i)}\in\mathbb{R}^{|\mathcal{D}_{k}^{(i)}|\times|\mathcal{D}_{k}^{(i)}|}$ for each client $k$ and task $i$, and a globally shared matrix $Q^{(i)}\in\mathbb{R}^{d\times d}$, where $d$ denotes the dimension of the feature. Each client encrypts its local feature matrix $X_{k}^{(i)}=M_{g}(\mathcal{D}_{k}^{(i)})$ via Inverse Singular Value Decomposition (ISVD):
(1) $X_{k}^{(i)^{\prime}}=P_{k}^{(i)}X_{k}^{(i)}Q^{(i)},$
@@
(2) $X^{(i)^{\prime}}=\text{concat}\{X_{k}^{(i)^{\prime}}\mid k=1,\ldots,K\},$
@@
(3) $X^{(i)^{\prime}}=U^{(i)^{\prime}}\Sigma^{(i)^{\prime}}V^{(i)^{\prime}\top},$
@@
(4) $U_{k}^{(i)}=\mathcal{I}_{k}(U^{(i)^{\prime}})\in\mathbb{R}^{|\mathcal{D}_{k}^{(i)}|\times n},$
where $\mathcal{I}_{k}(\cdot)$ denotes a row selection function that returns the indices corresponding to client $k$’s samples. To quantify the importance of local samples within the global latent space, client $k$ computes a leverage score (Drineas et al., 2012) for $j$-th sample of task $i$ as:
(5) $\tau_{k}^{i,j}=\|e_{i,j}^{\top}U_{k}^{(i)}\|_{2}^{2},$
@@
(6) $p_{k}^{i,j}=\frac{\tau_{k}^{i,j}}{\sum_{j^{\prime}=1}^{n_{k}^{i}}\tau_{k}^{i,j^{\prime}}}.$
@@
4.2. Task-aware Temperature Scaling (TTS)
@@
4.3. Training Strategy
The training strategy consists of two stages to progressively address the evolving challenges in federated class-incremental learning.
@@ 4.3. Training Strategy
(8) $\min_{\theta_{k}}\;\frac{1}{N}{\textstyle\sum_{i=1}^{N}}\mathrm{CE}(y_{i},\mathrm{Softmax}(f_{\theta_{k}}(x_{i}))),$
@@
5. Experiments
5.1. Experiment Settings
@@
Following existing studies (Zhang et al., 2023; Li et al., 2024), we conducted all experiments on three commonly used datasets, including CIFAR10 (Krizhevsky et al., 2009), CIFAR100 (Krizhevsky et al., 2009) and TinyImageNet (Le and Yang, 2015) to validate the effectiveness of the FedCBDR. We simulate heterogeneous data distributions across clients using the Dirichlet distribution with parameters $\beta=\{0.1,0.5,1.0\}$, where smaller values of $\beta$ correspond to higher level of data heterogeneity.
@@
Following prior studies (Zhang et al., 2023; Tran et al., 2024), we adopt Top-1 Accuracy as the evaluation metric, defined as $\text{Accuracy}=N_{\text{correct}}/N_{\text{total}}$, where $N_{\text{correct}}$ and $N_{\text{total}}$ denote the number of correct predictions and the total number of samples, respectively.
@@
Table 1. Performance comparison between FedCBDR and baselines across three datasets under varying levels of heterogeneity ($\beta$). CIFAR10 is divided into 3 tasks, CIFAR100 into 5 tasks, and TinyImageNet into 10 tasks. All methods were executed under three different random seeds, and both the mean and standard deviation of the results are reported. The best results are bolded.
@@
Table 2. Performance comparison between FedCBDR and baselines across three datasets under varying levels of heterogeneity ($\beta$). CIFAR10 is divided into 5 tasks, CIFAR100 into 10 tasks, and TinyImageNet into 20 tasks. All methods were executed under three different random seeds, and both the mean and standard deviation of the results are reported. The best results are bolded.
@@
5.2. Performance Comparison
To evaluate the effectiveness of the proposed FedCBDR, we compare it with six representative baseline methods: Finetune (Tran et al., 2024), FedEWC (Kirkpatrick et al., 2017), FedLwF (Li and Hoiem, 2017), TARGET (Zhang et al., 2023), LANDER (Tran et al., 2024), and Re-Fed (Li et al., 2024). As reported in Table 1 and Table 2, the results can be summarized as follows:
@@ 5.3. Ablation Study
  • Using the TTS module alone leads to consistent improvements over Finetune, highlighting its effectiveness in addressing intra-client class imbalance through temperature scaling. This contribution to better generalization is particularly evident under the more challenging ”5/10/20” task splitting scenario.
@@
Table 3. Ablation results under different levels of data heterogeneity and task splitting settings. “3/5/10” denotes CIFAR10 with 3 tasks, CIFAR100 with 5 tasks, and TinyImageNet with 10 tasks; “5/10/20” represents 5, 10, and 20 tasks respectively.
Task Splitting Method CIFAR10 CIFAR100 TinyImageNet
@@
5.4. Performance Evaluation of FedCBDR under Incremental Tasks
Figure 3. Performance comparison of all methods across varying task splits on CIFAR10 (3/5 tasks), CIFAR100 (5/10 tasks), and TinyImageNet (10/20 tasks) with $\beta=0.5$.

### v4

# Class-wise Balancing Data Replay for Federated Class-Incremental Learning

@@ Abstract
1 Introduction
@@
Federated learning (FL) is a distributed machine learning paradigm that enables collaborative training of a shared global model across multiple data sources FedSSA; hu2024fedmut; haozhao1; feng2024federated; fuless. It periodically performs parameter-level interaction between clients and the server instead of gathering clients’ data, which can enhance data privacy while leveraging the diversity of distributed data sources to build a more generalized global model liao2025privacy; qi2023cross; fu2025learn; hu2024fedcross; wang2024feddse; FedGH. This mechanism makes it widely applicable to various fields liang2025tta; feng2025scalable; qi2024attentive; zhong2025unlearning; zhong2022flee. Building upon this foundation, Federated Class-Incremental Learning (FCIL) extends FL by introducing dynamic data streams where clients sequentially encounter different task classes under non-independent and identically distributed data dong2022federated; wu2025federated; tran2024text; yang2024federated; fu2025beyond. However, this amplifies the inherent complexities of FL, as the global model must integrate heterogeneous and evolving knowledge from clients while mitigating catastrophic forgetting, despite having no or only limited access to historical data lu2024federated; chen2024general; gao2024fedprok.
@@
Figure 1: Motivation of the FedCBDR. Traditional data replay strategies typically focus on local information and, due to the lack of global awareness, often result in imbalanced class distributions during replay. FedCBDR aims to explore global information in a privacy-preserving manner and leverage it for sampling, which can alleviate the class imbalance problem.
@@
To address the challenge of catastrophic forgetting in FCIL, data replay has emerged as a promising strategy for retaining knowledge from previous tasks. Existing replay-based methods can be broadly categorized into two types: generative-based replay and exemplar-based replay. The former leverages generative models to synthesize representative samples from historical tasks chen2024general; babakniya2023data; wang2024data. Its core idea is to learn the data distribution of previous tasks and internalize knowledge in the form of model parameters, enabling the indirect reconstruction of prior knowledge through sample generation when needed tran2024text; zhang2023target. However, they often overlook the computational cost of training generative models and are inherently constrained by the quality and fidelity of the synthesized data chen2024general; liang2024diffusion; yoo2024federated. In contrast, exemplar-based replay methods directly store real samples from previous tasks, avoiding the complexity of generative processes while leveraging high-quality raw data to ensure robust retention of prior knowledge dong2022federated; liang2024diffusion; li2024towards; li2025re. These methods rely on a limited set of historical samples to maintain the decision boundaries of previously learned task classes. However, due to the lack of a global perspective on data distribution across clients, these methods are prone to class-level imbalance in replayed samples, which undermines the model’s ability to retain prior knowledge li2024towards; li2025re.
@@
2 Related Work
2.1 Exemplar-based Replay Methods
@@
In FCIL, exemplar-based replay methods aim to mitigate catastrophic forgetting by storing and replaying a subset of samples from previous tasks. They typically maintain a small exemplar buffer on each client, which is used during training alongside new task data to preserve knowledge of previously learned classes dong2022federated; li2024towards; li2025re; rebuffi2017icarl; 10323204; li2024sr. For example, GLFC alleviates forgetting in FCIL by leveraging local exemplar buffers for rehearsal, while introducing class-aware gradient compensation and prototype-guided global coordination to jointly address local and global forgetting dong2022federated. Moreover, Re-Fed introduces a Personalized Informative Model to strategically identify and replay task-relevant local samples, enhancing the efficiency of buffer usage and further reducing forgetting in heterogeneous client environments li2024towards. However, the lack of global insight in local sample selection often results in class imbalance, while the long-tailed distribution between replayed and current data is frequently overlooked, degrading the effectiveness of data replay li2024towards; li2025re.
2.2 Generative-based Replay Methods
@@
Generative replay methods aim to reconstruct the samples of past tasks through techniques such as generative modeling tran2024text; chen2024general; zhang2023target; nguyen2024overcoming; qi2023better, which enables the model to revisit historical knowledge to mitigate catastrophic forgetting. Following this line of thought, TARGET generates pseudo features through a globally pre-trained encoder and performs knowledge distillation by aligning the current model’s predictions with those of a frozen global model zhang2023target; LANDER utilizes pre-trained semantic text embeddings as anchors to synthesize meaningful pseudo samples, and distills knowledge by aligning the model’s predictions with class prototypes derived from textual descriptions tran2024text. However, these methods are typically limited by the high computational cost of training generative models and the suboptimal performance caused by low-fidelity pseudo samples tran2024text; zhang2023target.
2.3 Knowledge Distillation-based Methods
@@
Knowledge distillation-based methods generally follow two paradigms. The first focus om aligning the output predictions of the current model with those of previous models, which aims to preserve task-specific decision boundaries li2017learning; tan2024fl; xu2025self; psaltis2023fedrcil; chen2025knowledge; zhong2025sacfl; xu2024distribution; dong2023federated_FISS; xu2025long. The second estimates the importance of model parameters for previously learned tasks and performs regularization to prevent forgetting kirkpatrick2017overcoming; yu2024overcoming. Both approaches avoid storing raw data but are prone to knowledge degradation over time, especially as the number of tasks increases li2017learning; kirkpatrick2017overcoming.
3 Preliminaries
@@
4 Class-wise Balancing Data Replay for Federated Class-Incremental Learning
This section presents an effective active data selection method for FCIL, which aims to explore global data distribution to balance class-wise sampling. Moreover, it leverages temperature scaling to adjust the logits, which can alleviate the imbalance between samples from previously learned and newly introduced tasks. Figure 2 and Algorithm 1 illustrates the framework of the proposed FedCBDR.
4.1 Global-perspective Data Replay (GDR)
@@
Figure 2: Illustration of the FedCBDR framework. It first trains local models using samples from current and previous tasks. After a fixed number of communication rounds, each client extracts local sample features using the global model and applies inverse singular value decomposition (ISVD) to obtain pseudo features. The server then aggregates both local models and pseudo features, performs SVD on the features, and selects representative samples based on leverage scores. The corresponding sample indices are sent back to the clients for balanced replay.
Inspired by Singular Value Decomposition (SVD) chai2024efficient; an2022numerical, we first generate a set of random orthogonal matrices: a client-specific matrix $P_{k}^{(i)}\in\mathbb{R}^{|\mathcal{D}_{k}^{(i)}|\times|\mathcal{D}_{k}^{(i)}|}$ for each client $k$ and task $i$, and a globally shared matrix $Q^{(i)}\in\mathbb{R}^{d\times d}$, where $d$ denotes the dimension of the feature. Each client encrypts its local feature matrix $X_{k}^{(i)}=M_{g}(\mathcal{D}_{k}^{(i)})$ via Inverse Singular Value Decomposition (ISVD):
$X_{k}^{(i)^{\prime}}=P_{k}^{(i)}X_{k}^{(i)}Q^{(i)},$ (1)
@@
$X^{(i)^{\prime}}=\text{concat}\{X_{k}^{(i)^{\prime}}\mid k=1,\ldots,K\},$ (2)
@@
$X^{(i)^{\prime}}=U^{(i)^{\prime}}\Sigma^{(i)^{\prime}}V^{(i)^{\prime}\top},$ (3)
@@
$U_{k}^{(i)}=\mathcal{I}_{k}(U^{(i)^{\prime}})\in\mathbb{R}^{|\mathcal{D}_{k}^{(i)}|\times n},$ (4)
where $\mathcal{I}_{k}(\cdot)$ denotes a row selection function that returns the indices corresponding to client $k$’s samples. To quantify the importance of local samples within the global latent space, client $k$ computes a leverage score drineas2012fast; pan2016efficient; yu2011dynamic for $j$-th sample of task $i$ as:
$\tau_{k}^{i,j}=\|e_{i,j}^{\top}U_{k}^{(i)}\|_{2}^{2},$ (5)
@@
$p_{k}^{i,j}=\frac{\tau_{k}^{i,j}}{\sum_{j^{\prime}=1}^{n_{k}^{i}}\tau_{k}^{i,j^{\prime}}}.$ (6)
@@
Algorithm 1 FedCBDR
1:Initialize: $R$: number of communication rounds; $K$: number of clients; $t$: number of tasks; $\theta_{g}$: global model parameters; $\mathcal{B}_{k}^{pre}$: replay buffer for historical tasks on client $k$; $\mathcal{D}_{k}^{s}$: local data of task $s$ on client $k$.
2:for each task $s=1$ to $t$ do
3:  for each communication round $r=1$ to $R$ do
4:   for each client $k=1$ to $K$ do
5:    Initialize local model parameters: $\theta_{k}\leftarrow\theta_{g}$
6:    if $s==1$ then
7:     Sample a mini-batch $\zeta$ from $\mathcal{D}_{k}^{(1)}$, and update $\theta_{k}$ using Eq. 8.
8:    else
9:     Store the historical task data corresponding to globally sampled IDs into $\mathcal{B}_{k}^{pre}$.
10:     Sample a mini-batch $\zeta$ from $\mathcal{D}_{k}^{(s)}\cup\mathcal{B}_{k}^{pre}$, and update $\theta_{k}$ using Eq. 9.
11:     Compute pseudo-features based on Eq. 1, and upload them to the server.
12:    end if
13:   end for
14:   if $r<R$ then
15:    Aggregate local model parameters across clients.
16:   else
17:    Aggregate model parameters and pseudo-features from all clients using Eq. 2.
18:    Perform Global Sampling based on Eqs. 36, and send the selected sample IDs back to the corresponding clients.
19:   end if
20:  end for
21:end for
22:// Global Sampling Procedure
23: Form the global feature pool $X^{(i)}$ by aggregating all pseudo-features via Eq. (2).
24: Perform singular value decomposition (SVD) using Eq. 3 to extract key attributes.
25: Compute leverage scores for each client’s samples using Eqs. 45, and normalize globally using Eq. 6.
26: Perform sampling and adjust the probabilities of the selected samples accordingly.
4.2 Task-aware Temperature Scaling (TTS)
@@
4.3 Training Strategy
The training strategy consists of two stages to progressively address the evolving challenges in federated class-incremental learning. Algorithm 1 presents the pipeline of the FedCBDR.
@@ 4.3 Training Strategy
$\min_{\theta_{k}}\;\frac{1}{N}{\textstyle\sum_{i=1}^{N}}\mathrm{CE}(y_{i},\mathrm{Softmax}(f_{\theta_{k}}(x_{i}))),$ (8)
@@
5 Experiments
Table 1: Statistics of the datasets used in experiments.
Datasets #Class #Training #Testing Image Size Federated Settings
Clients Tasks Heterogeneity
CIFAR10 10 50,000 10,000 32 $\times$ 32 5/10 3/5 0.5/1.0
CIFAR100 100 50,000 10,000 32 $\times$ 32 5/10 5/10 0.1/0.5/1.0
TinyImageNet 200 100,000 10,000 64 $\times$ 64 5/10 10/20 0.1/0.5/1.0
5.1 Experiment Settings
@@
Following existing studies zhang2023target; li2024towards, we conducted all experiments on three commonly used datasets, including CIFAR10 krizhevsky2009learning; qi2022clustering, CIFAR100 krizhevsky2009learning; qi2022clustering and TinyImageNet le2015tiny to validate the effectiveness of the FedCBDR. We simulate heterogeneous data distributions across clients using the Dirichlet distribution with parameters $\beta=\{0.1,0.5,1.0\}$, where smaller values of $\beta$ correspond to higher level of data heterogeneity. The statistical details are presented in the Table 1.
@@
Following prior studies tran2024text; chen2024flexfl; wang2024fednlr; qi2024cross, we adopt Top-1 Accuracy as the evaluation metric, defined as $\text{Accuracy}=N_{\text{correct}}/N_{\text{total}}$, where $N_{\text{correct}}$ and $N_{\text{total}}$ denote the number of correct predictions and the total number of samples, respectively.
@@
Table 2: Performance comparison between FedCBDR and baselines across three datasets under varying levels of heterogeneity ($\beta$). CIFAR10 is divided into 3 tasks, CIFAR100 into 5 tasks, and TinyImageNet into 10 tasks. All methods were executed under three different random seeds, and both the mean and standard deviation of the results are reported. The best results are bolded.
@@
Table 3: Performance comparison between FedCBDR and baselines across three datasets under varying levels of heterogeneity ($\beta$). CIFAR10 is divided into 5 tasks, CIFAR100 into 10 tasks, and TinyImageNet into 20 tasks. All methods were executed under three different random seeds, and both the mean and standard deviation of the results are reported. The best results are bolded.
@@
5.2 Performance Comparison
To evaluate the effectiveness of the proposed FedCBDR, we compare it with six representative baseline methods: Finetune tran2024text, FedEWC kirkpatrick2017overcoming, FedLwF li2017learning, TARGET zhang2023target, LANDER tran2024text, and Re-Fed li2024towards. As reported in Table 2 and Table 3, the results can be summarized as follows:
@@ 5.3 Ablation Study
  • Using the TTS module alone leads to consistent improvements over Finetune, highlighting its effectiveness in addressing intra-client class imbalance through temperature scaling. This contribution to better generalization is particularly evident under the more challenging "5/10/20" task splitting scenario.
@@
Table 4: Ablation results under different levels of data heterogeneity and task splitting settings. “3/5/10” denotes CIFAR10 with 3 tasks, CIFAR100 with 5 tasks, and TinyImageNet with 10 tasks; “5/10/20” represents 5, 10, and 20 tasks respectively.
Task
Splitting
Method CIFAR10 CIFAR100 TinyImageNet
@@
5.4 Performance Evaluation of FedCBDR under Incremental Tasks
Figure 3: Performance comparison of all methods across varying task splits on CIFAR10 (3/5 tasks), CIFAR100 (5/10 tasks), and TinyImageNet (10/20 tasks) with $\beta=0.5$.